OpenStudy (anonymous):

Find the length of the arc on a circle of radius 9 inches that is intercepted by a central angle of 60 degrees

jimthompson5910 (jim_thompson5910):

what is the circumference of the circle?

OpenStudy (anonymous):

360

jimthompson5910 (jim_thompson5910):

that's how many degrees are in a circle

jimthompson5910 (jim_thompson5910):

I'm asking about the perimeter around the circle

OpenStudy (jdoe0001):

\(\bf\huge arc = \frac{r\theta\pi}{180}\)

OpenStudy (anonymous):

\[\theta/360\times \pi \times2r\]

OpenStudy (skullpatrol):



OpenStudy (anonymous):

thus u'l have something like 60/360\[\times \pi \times2\times9=l\]

OpenStudy (anonymous):

and your final answer will be 9.425

OpenStudy (anonymous):

alright so is it 3pi?

OpenStudy (anonymous):

ohhh okay, thanks (all of you) for the help!!! i get it